Telescope Pointing Errors Update

We started the telescope on a mapping run last night. The first ~20 points looked very good, but then the dome stopped responding to commands to rotate. The dome shutter could be opened and closed, but the dome could not be rotated
At this point we were forced to abandon the mapping run for the night, leaving us wondering what we would find wrong with the dome when we returned to the site in the morning. There are many possible points of failure in the dome system, many of which would result in the extension of the maintenance trip, and possibly having new parts shipped from America.

The most puzzling part of the dome failure is that we had already completed thorough maintenance checks and some small repair work on the dome earlier in the trip... only the return to the telescope today would reveal our fate.

On our arrival at the observatory we quickly discovered that the resettable dome Az motor fuse had blown at some point in the night, but there were no immediate explanations of why. We reset the fuse and tested the rotation of the dome. Everything appeared to be operating normally until we heard an unpleasant noise and spotted that there was a cable hanging down from the upper ring of the dome which shouldn't have been there.

We found that the cable for the upper dome paddle (that can open, stop and shut the dome) had been ripped from its connector on the side of the upper dome control box. We are speculating that this loose cable caught in the dome Az mechanism and caused the dome to stall, triggering the fuse failure. This is very strange, as the control paddle, control box and cables have all remained in the same position untouched for a number of years... but I have now repaired the cable and routed the cable differently to try to avoid a repeat of this problem.

We are now attempting another alignment run, but this time our attempts may fall victim to the weather.

No comments:

Post a Comment